### v2

@@ Abstract
We continue the study of 2D gravity — “matrix model” duality on the example of $(2,2p+1)$ minimal string. We propose a reformulation of the duality, related to a more conventional one by “$x-y$ swap” in the language of topological recursion. This formulation elucidates some conceptual and technical difficulties in the dictionary of the duality and relation to other examples. In particular, it allows to circumvent the necessity to use “resonance transformations”, that were previously introduced to match the worldsheet and “matrix model” correlators, and the expressions for minimal string amplitudes in this approach are reminiscent of the ones obtained recently for “complex Liouville string” theory. Using this new approach, we also formulate a conjecture on how one can compute amplitudes with operators other than tachyons in the dual theory.
@@ 1 Introduction
An important property characterizing the dual description is the recurrent relations for correlators, known as topological recursion (TR), which in matrix model roughly corresponds $1/N$ expansion of loop equations. Topological recursion has been thoroughly studied on its own in the mathematical literature and is known to have a lot of interesting properties. For example, TR data can be expressed in terms of tautological intersection theory on the moduli space of stable curves Eynard:2011sc , a connection that bridges the gap between TR and worldsheet interpretation in terms of “Euler characteristic of bundle of conformal blocks” for VMS.
Another one of these properties, studied in a different language in older works Kharchev:1993as ; KawaiFuk and recently investigated in more generality Alexandrov:2022ydc , is the transformation of TR data under the so-called “$x-y$ swap”. A relevant example of a system of correlators satisfying TR to have in mind are resolvent correlators in (e.g. hermitian) two-matrix models, where one integrates over matrices $X$ ad $Y$ with the weight of the form $\exp(-V(X,Y))$, $V(X,Y)=\text{tr\,}\left(V_{1}(X)+V_{2}(Y)+XY\right)$. In this case, “$x-y$ swap” transformation relates the correlators of resolvents that depend on $X$ and $Y$ in this model. Putative dual of minimal string can be described in this way (more precisely, as a “double-scaling limit” of such models).
For $(2,2p+1)$ series, it reduces to a one-matrix model, if we are only interested in the observables that depend on $X$ (one can think that in thhe two-matrix model $V_{2}(Y)$ is quadratic and $Y$ can be integrated out). This is the language in which the duality is usually formulated. However, we would like to argue that for comparison with the worldsheet it might be more natural to think of “$Y$-dependent” observables instead. The main proposition of the current paper is that the map between MS worldsheet correlators and matrix model one, previously described in somewhat obscure terms, can be much more easily formulated using TR after “$x-y$ swap”.
@@ 2.1 Definitions and notations
Primary fields of Liouville theory exist for any conformal dimension $\Delta={\overline{\Delta}}$. We will denote them in several different ways: $V_{a}$ is an operator of dimension $\Delta=a(Q-a)$, with $a$ identified with $Q-a$; alternatively, one can parametrize it with Liouville momentum $P$: $\Delta=\frac{Q^{2}}{4}+P^{2}$, then $P$ and $-P$ are identified. The spectrum of Liouville theory, i.e. fields that appear in the OPE, have $P$ real. In minimal string, Liouville fields with $a=\frac{(1-m)b^{-1}+(1-n)b}{2}\equiv a_{m,n}$, or $P=i\frac{mb^{-1}+nb}{2}\equiv P_{m,n}$ frequently appear. These fields will be denoted also by $V_{m,n}$. For positive integer $m,n$ these fields are degenerate on level $mn$: a null descendant of the form $D_{m,n}^{(L)}V_{m,n}=0$ exists in the corresponding Verma module.
@@
Finally, the ghosts are a free fermionic (with anticommuting fields) CFT that has two holomorphic fundamental fields $B,C$ with conformal dimensions $(2,0)$ and $(-1,0)$ respectively and two antiholomorphic ${\overline{B}},{\overline{C}}$. The central charge of ghost CFT is $-26$. In the context of string theory, on nontrivial closed Riemann surface (e.g. torus) they are to be quantized with periodic boundary conditions along the cycles (see e.g. polchinski_1998 ).
@@ 2.2 Operators: tachyons and ground ring
Tachyon amplitudes should be symmetric in the parameters of the correlator. In particular, for genus zero, when integrating over moduli space is equivalent to integration over $n-3$ cross-ratios, it should not depend on which $n-3$ fields are integrated.
@@
where $H$ is a polynomial of degree $n-1$ of ghost and Virasoro modes $L_{-k},M_{-k}$ (in Liouville and minimal model sector respectively). They are built from Virasoro descendants; the reason why they are, nevertheless, BRST-invariant, is because when acting with $\mathcal{Q}$ the result is a linear combination of null descendants $D^{(L)}_{m,n}V_{m,n}$ and the corresponding one in minimal model $D^{(M)}_{m,n}\Phi_{m,n}$(which we put to zero). This is the defining relation that allows to find the explicit form of $H_{1,n}$ case by case imbimbo1992 (although no general formula is known). The main example is $O_{1,2}$ with
@@
$:2$” denotes summation with step $2$ here and in what follows. The last two equations are valid in cohomology, i.e. up to $\mathcal{Q}$-exact terms; alternatively, their RHS gives a leading (independent of $x$) term in the OPE $O(x)O(0)$ and $O(x)W(0)$ respectively. The coefficients $G$ and $A$ in fact can be put to one if one considers renormalized operators $\mathcal{O}_{1,n}=\frac{O_{1,n}}{\Lambda_{1,n}},\,\mathcal{W}_{1,n}=\frac{W_{1,n}}{\mathcal{N}(a_{1,-n})}$ Belavin:2005jy .
@@ 2.3 Available results from higher equations of motion
Additional terms come from non-diagonal action of $L_{0}$ on logarithmic operators. But after acting with an antiholomorphic operator (¯$({\overline{\partial}}{\overline{H}}_{1,n}-{\overline{\mathcal{Q}R}}_{1,n})$, it disappears, since $\mathcal{{\overline{D}}}_{1,n}\Theta_{1,n}=0$. In total we are left with
[…]
Thus, HEM allow to represent the integrated tachyon as a total derivative ¯$\partial{\overline{\partial}}O^{\prime}_{1,n}$ up to BRST-exact terms. If all operators other than this tachyon are BRST-closed, at least formally BRST-exact contributions can be omitted; then, the integral reduces to a sum of boundary terms in the vicinity of other insertions. These boundary terms can be calculated, because only a finite number of contributions in the Liouville OPE are relevant for this. Namely, one can write
@@ 2.4 Discontinuity analysis
×$\times\sum\limits_{\hat{P}}C^{M}(iP_{1},iP_{2},\hat{P})C^{M}(iP_{3},iP_{4},\hat{P})|\mathcal{F}^{M,ib}(\hat{P},iP_{i}|z)|^{2}(z{\overline{z}})^{P^{2}+\hat{P}^{2}-1}$
@@
To analytically continue to this regime, we can first perform an integral over $z$. Starting with a discrete term at intermediate momentum $P^{*}$ s.t. Re $\text{Re\,}P^{*2}>0$, if conformal blocks are expanded to leading order, integral over $z$ over the vicinity of $z=0$ gives
@@ 3.1 MS spectral curve
We use an unconventional way to enumerate them; a usual convention is $t_{k}=\frac{2p-2k+1}{2}t^{usual}_{2p-2k+1}$. We will only consider $t_{k}$ for $k=1,...,p$ for now. One can check that even times $t_{2k}$ in this set are zero. The derivatives themselves are given by Alexandrov:2023jcj
@@ 3.3 $p$-deformed volumes
4 $x-y$ swapped spectral curve and correlation numbers
@@
We now proceed to one of the main statements of this work: an alternative prescription to calculate $A^{g}_{n}$, which we believe to be able to (at least partially) solve all the caveats listed above. The proposal is as follows: the $x-y$ dual spectral curve with
$\begin{cases}\check{x}=2u_{0}^{\frac{2p+1}{2}}T_{2p+1}(z)\\ \check{y}=2u_{0}\,T_{2}(z)\\ \end{cases}$ (31)
@@ 4.2 Explicit expressions for tachyon correlators and checks at low $g,n$
$\sum\limits_{m=1}^{2p}\frac{(-1)^{m}}{2(2p+1)^{2}}\frac{\sin^{2}\frac{\pi m}{2p+1}}{\cos\frac{\pi m}{2p+1}}\frac{\prod\limits_{i=0}^{2}\sin\frac{\pi m(2(p-k_{i})+1)}{(2p+1)}}{\sin^{3}\frac{\pi m}{2p+1}}=\frac{1}{2(2p+1)}\cdot b^{2}\sum\limits_{m=1}^{2p}\frac{\prod\limits_{i=0}^{2}\sin\frac{2\pi mk_{i}}{(2p+1)}}{\sin\pi mb^{2}}$ (39)
@@
and $\circ$ and $\bullet$ in the first term denote two arguments other than 00 and $i$ from the set $\{0,1,2,3\}$. Then one can perform a “Chebyshev transform” (32). It can be brought to a simpler form if we represent $(m,n)$ dependent kernels in the expressions above, appropriately symmetrized with respect to $m\leftrightarrow 2p+1-m$, as sums of the form $\sum\limits_{k=1}^{2p}\sin(\pi kmb^{2})\sin(\pi knb^{2})f(k)$. This rewriting allows to associate a Verlinde-like factor to each summation variable. For our kernels it turns out that $f(k)$ are simple enough — polynomial in $|p+1/2-k|$. E.g. [^4] Note that the identity below is only valid for $m,n\in 1\dots 2p$. It is not true if we consider both sides as functions of real variables $m,n$, which complicates proof of such identities.
@@
$2b^{2}\left(k_{1}(k_{1}+p+\frac{1}{2})-\sum\limits_{i=2,3,4}\sum\limits_{m=k_{i}-k_{1}+1:2}^{k_{i}+k_{1}-1}|p+\frac{1}{2}-m|\right)$ (48)
This is the answer obtained for 4-point correlator using “higher equations of motion” (compare with (15), up to normalization, or with the expression in belzam2009 ), which is valid precisely under the assumption $\mathcal{N}^{(0)}_{k_{1}\dots k_{4}}=k_{1}$. Finally, we can perform some explicit numerical checks; figure 1 illustrates that or formula coincides with the result in belzam2009 , in particular, reduces to $p$-deformed volume in the appropriate domain.
@@
  • this is a connected graph with $n$ external “legs”, vertices and the edges connecting some of the vertices. Legs correspond to marked points, vertices — to smooth components of degenerate surface and edges to the nodal points where degeneration happens.
  • With each vertex $v$, a number $g_{v}\in\mathbb{Z}_{\geq 0}$ is associated which is a genus of the corresponding smooth component. For each $v$ a stability condition holds: $02g_{v}-2+n_{v}>$ 0, where $n_{v}$ is the valence of $v$ including edges and legs.
@@
However, the next examples that are presented in the appendix C$(g,n)=(0,5)$ and $(g,n)=(1,2)$, although exhibit the expected structure as sum over “stable graphs”, contain additional contributions that do not follow from the diagrammatic technique above.
@@
×$\times\left(\prod\limits_{e}\left(\frac{-\partial}{4b^{2}\partial(P_{1,-k_{e}}^{2})}\right)^{n_{e}-1}\prod\limits_{v}\frac{V_{g_{v},n_{v}}^{b}(4\pi bP_{1,-\vec{k}_{v}})}{(2\pi^{2})^{3g_{v}+n_{v}-3}}\right)\Bigg{|}_{P_{e}^{2}=0}$ (57)
The sum over $\Gamma$ includes all stable graphs of genus $g$ with $n$ legs with external markings $k_{1}\dots k_{n}$. For each edge in the graph, a marking $k_{e}$ and a natural number $n_{e}$ are associated and the sum is over all these assignments. It is simpler to write this answer as a sum of tautological intersection numbers on ¯${\overline{\mathcal{M}}}_{\Gamma}=\prod\limits_{v}{\overline{\mathcal{M}}}_{g_{v},n_{v}}$ — moduli space of stable curves with the degenerations specified by stable graph $\Gamma$, as follows:
@@
×$\times\,e^{\sum\limits_{i=1}^{n}(2bP_{1,-k_{i}})^{2}\psi_{i}}\times\sum\limits_{\vec{n_{e}};\,n_{e}=1}^{\infty}\prod\limits_{e}\frac{4^{n_{e}}B_{2n_{e}}(|bP_{1,-k_{e}}|)(-\psi_{e\bullet}-\psi_{e\circ})^{n_{e}-1}}{n_{e}!}$ (58)
In product over $v$ it is understood that each $\kappa$-class is an element in cohomology $H^{\bullet}({\overline{\mathcal{M}}}_{g_{v},n_{v}})$. We will not give precise definitions of $\kappa$ and $\psi$-classes and refer to expositions in Zvonkine2012AnIT ; Collier:2023cyw .
@@
It formally reduces to a series in residues in the right half-plane (at each order in $u$ the series for Hurwitz zeta-function $\zeta(-2n+1,bP)$ is obtained, which reduces to Bernoulli polynomials). This series takes the form
@@
The sine factors appear from Verlinde representation of fusion numbers appearing at the vertices $v_{1},v_{2}$. First, note that this is just a “discrete version” of the “propagator” in CLS Collier:2024matrixmodel , where the integral is replaced by a sum. Second, note that $P_{1,-n}$ in the total sum go through all the values in Lian-Zuckerman set, enumerating physical operators. The restriction $n\text{ mod }(2p+1)\neq 0$ is in fact not required, since such terms automatically disappear because of sine factors from the fusion rules.
If we substitute the series (59) into (58), we can perform integral over ¯${\overline{\mathcal{M}}}_{\Gamma}$ explicitly in terms of “quantum volumes”; recall that their definition in our normalization is

### v3

@@ Abstract
We continue the study of 2D gravity — “matrix model” duality on the example of $(2,2p+1)$ minimal string. We propose a reformulation of the duality, related to a more conventional one by “xyx-y swap” in the language of topological recursion. This formulation elucidates some conceptual and technical difficulties in the dictionary of the duality and relation to other examples. In particular, it allows to circumvent the necessity to use “resonance transformations”, that were previously introduced to match the worldsheet and “matrix model” correlators, and the expressions for minimal string amplitudes in this approach are reminiscent of the ones obtained recently for “complex Liouville string” theory. Using this new approach, we also formulate a conjecture on how one can compute amplitudes with operators other than tachyons in the dual theory.
@@ 1 Introduction
An important property characterizing the dual description is the recurrent relations for correlators, known as topological recursion (TR), which in matrix model roughly corresponds to $1/N$ expansion of loop equations. Topological recursion has been thoroughly studied on its own in the mathematical literature and is known to have a lot of interesting properties. For example, TR data can be expressed in terms of tautological intersection theory on the moduli space of stable curves Eynard:2011sc , a connection that bridges the gap between TR and worldsheet interpretation in terms of “Euler characteristic of bundle of conformal blocks” for VMS.
Another one of these properties, studied in a different language in older works Kharchev:1993as ; KawaiFuk and recently investigated in more generality Alexandrov:2022ydc , is the transformation of TR data under the so-called “xyx-y swap”. A relevant example of a system of correlators satisfying TR to have in mind are resolvent correlators in (e.g. hermitian) two-matrix models, where one integrates over matrices $X$ ad $Y$ with the weight of the form $\exp(-V(X,Y))$, $V(X,Y)=\text{tr\,}\left(V_{1}(X)+V_{2}(Y)+XY\right)$. In this case, “xyx-y swap” transformation relates the correlators of resolvents that depend on $X$ and $Y$ in this model. Putative dual of minimal string can be described in this way (more precisely, as a “double-scaling limit” of such models).
For $(2,2p+1)$ series, it reduces to a one-matrix model, if we are only interested in the observables that depend on $X$ (one can think that in thhe two-matrix model $V_{2}(Y)$ is quadratic and $Y$ can be integrated out). This is the language in which the duality is usually formulated. However, we would like to argue that for comparison with the worldsheet it might be more natural to think of “$Y$-dependent” observables instead. The main proposition of the current paper is that the map between MS worldsheet correlators and matrix model one, previously described in somewhat obscure terms, can be much more easily formulated using TR after “xyx-y swap”.
@@ 2.1 Definitions and notations
Primary fields of Liouville theory exist for any conformal dimension $\Delta={\overline{\Delta}}$. We will denote them in several different ways: $V_{a}$ is an operator of dimension $\Delta=a(Q-a)$, with $a$ identified with QaQ-a; alternatively, one can parametrize it with Liouville momentum $P$: $\Delta=\frac{Q^{2}}{4}+P^{2}$, then $P$ and P-P are identified. The spectrum of Liouville theory, i.e. fields that appear in the OPE, have $P$ real. In minimal string, Liouville fields with $a=\frac{(1-m)b^{-1}+(1-n)b}{2}\equiv a_{m,n}$, or $P=i\frac{mb^{-1}+nb}{2}\equiv P_{m,n}$ frequently appear. These fields will be denoted also by $V_{m,n}$. For positive integer $m,n$ these fields are degenerate on level $mn$: a null descendant of the form $D_{m,n}^{(L)}V_{m,n}=0$ exists in the corresponding Verma module.
@@
Finally, the ghosts are a free fermionic (with anticommuting fields) CFT that has two holomorphic fundamental fields $B,C$ with conformal dimensions $(2,0)$ and (1,0)(-1,0) respectively and two antiholomorphic ${\overline{B}},{\overline{C}}$. The central charge of ghost CFT is 26-26. In the context of string theory, on nontrivial closed Riemann surface (e.g. torus) they are to be quantized with periodic boundary conditions along the cycles (see e.g. polchinski_1998 ).
@@ 2.2 Operators: tachyons and ground ring
Tachyon amplitudes should be symmetric in the parameters of the correlator. In particular, for genus zero, when integrating over moduli space is equivalent to integration over n3n-3 cross-ratios, it should not depend on which n3n-3 fields are integrated.
@@
where $H$ is a polynomial of degree n1n-1 of ghost and Virasoro modes $L_{-k},M_{-k}$ (in Liouville and minimal model sector respectively). They are built from Virasoro descendants; the reason why they are, nevertheless, BRST-invariant, is because when acting with $\mathcal{Q}$ the result is a linear combination of null descendants $D^{(L)}_{m,n}V_{m,n}$ and the corresponding one in minimal model $D^{(M)}_{m,n}\Phi_{m,n}$(which we put to zero). This is the defining relation that allows to find the explicit form of $H_{1,n}$ case by case imbimbo1992 (although no general formula is known). The main example is $O_{1,2}$ with
@@
:2:2” denotes summation with step $2$ here and in what follows. The last two equations are valid in cohomology, i.e. up to $\mathcal{Q}$-exact terms; alternatively, their RHS gives a leading (independent of $x$) term in the OPE $O(x)O(0)$ and $O(x)W(0)$ respectively. The coefficients $G$ and $A$ in fact can be put to one if one considers renormalized operators $\mathcal{O}_{1,n}=\frac{O_{1,n}}{\Lambda_{1,n}},\,\mathcal{W}_{1,n}=\frac{W_{1,n}}{\mathcal{N}(a_{1,-n})}$ Belavin:2005jy .
@@ 2.3 Available results from higher equations of motion
Additional terms come from non-diagonal action of $L_{0}$ on logarithmic operators. But after acting with an antiholomorphic operator $({\overline{\partial}}{\overline{H}}_{1,n}-{\overline{\mathcal{Q}R}}_{1,n})$, it disappears, since $\mathcal{{\overline{D}}}_{1,n}\Theta_{1,n}=0$. In total we are left with
$(\partial H_{1,n}-\mathcal{Q}R_{1,n})({\overline{\partial}}{\overline{H}}_{1,n}-{\overline{\mathcal{Q}R}}_{1,n})\Theta^{\prime}_{1,n}=\mathcal{D}_{1,n}{\overline{\mathcal{D}}}_{1,n}\Theta^{\prime}_{1,n}=D^{(L)}_{1,n}{\overline{D}}^{(L)}_{1,n}\Theta^{\prime}_{1,n}\overset{\text{HEM}}{=}B_{1,n}U_{1,n}$ (11)
Thus, HEM allow to represent the integrated tachyon as a total derivative $\partial{\overline{\partial}}O^{\prime}_{1,n}$ up to BRST-exact terms. If all operators other than this tachyon are BRST-closed, at least formally BRST-exact contributions can be omitted; then, the integral reduces to a sum of boundary terms in the vicinity of other insertions. These boundary terms can be calculated, because only a finite number of contributions in the Liouville OPE are relevant for this. Namely, one can write
@@ 2.4 Discontinuity analysis
$\times\sum\limits_{\hat{P}}C^{M}(iP_{1},iP_{2},\hat{P})C^{M}(iP_{3},iP_{4},\hat{P})|\mathcal{F}^{M,ib}(\hat{P},iP_{i}|z)|^{2}(z{\overline{z}})^{P^{2}+\hat{P}^{2}-1}$
@@
To analytically continue to this regime, we can first perform an integral over $z$. Starting with a discrete term at intermediate momentum $P^{*}$ s.t. $\text{Re\,}P^{*2}>0$, if conformal blocks are expanded to leading order, integral over $z$ over the vicinity of $z=0$ gives
@@ 3.1 MS spectral curve
We use an unconventional way to enumerate them; a usual convention is $t_{k}=\frac{2p-2k+1}{2}t^{usual}_{2p-2k+1}$. We will only consider $t_{k}$ for k=1,,pk=1,...,p for now. One can check that even times $t_{2k}$ in this set are zero. The derivatives themselves are given by Alexandrov:2023jcj
@@ 3.3 $p$-deformed volumes
4 xyx-y swapped spectral curve and correlation numbers
@@
We now proceed to one of the main statements of this work: an alternative prescription to calculate $A^{g}_{n}$, which we believe to be able to (at least partially) solve all the caveats listed above. The proposal is as follows: the xyx-y dual spectral curve with
[…]
@@ 4.2 Explicit expressions for tachyon correlators and checks at low $g,n$
$\check{A}^{0}_{3}(\vec{k})=\sum\limits_{m=1}^{2p}\frac{(-1)^{m}}{2(2p+1)^{2}}\frac{\sin^{2}\frac{\pi m}{2p+1}}{\cos\frac{\pi m}{2p+1}}\frac{\prod\limits_{i=0}^{2}\sin\frac{\pi m(2(p-k_{i})+1)}{(2p+1)}}{\sin^{3}\frac{\pi m}{2p+1}}=\frac{1}{2(2p+1)}\cdot b^{2}\sum\limits_{m=1}^{2p}\frac{\prod\limits_{i=0}^{2}\sin\frac{2\pi mk_{i}}{(2p+1)}}{\sin\pi mb^{2}}$ (39)
@@
and $\circ$ and $\bullet$ in the first term denote two arguments other than 0 and $i$ from the set $\{0,1,2,3\}$. Then one can perform a “Chebyshev transform” (32). It can be brought to a simpler form if we represent $(m,n)$ dependent kernels in the expressions above, appropriately symmetrized with respect to $m\leftrightarrow 2p+1-m$, as sums of the form $\sum\limits_{k=1}^{2p}\sin(\pi kmb^{2})\sin(\pi knb^{2})f(k)$. This rewriting allows to associate a Verlinde-like factor to each summation variable. For our kernels it turns out that $f(k)$ are simple enough — polynomial in |p+1/2k||p+1/2-k|. E.g. [^4] Note that the identity below is only valid for $m,n\in 1\dots 2p$. It is not true if we consider both sides as functions of real variables $m,n$, which complicates proof of such identities.
@@
$\check{A}^{0}_{4}(\vec{k})=\frac{b^{4}}{4}\left(k_{1}(k_{1}+p+\frac{1}{2})-\sum\limits_{i=2,3,4}\sum\limits_{m=k_{i}-k_{1}+1:2}^{k_{i}+k_{1}-1}|p+\frac{1}{2}-m|\right)$ (48)
This is the answer obtained for 4-point correlator using “higher equations of motion” (compare with (15), up to normalization, or with the expression in belzam2009 ), which is valid precisely under the assumption $\mathcal{N}^{(0)}_{k_{1}\dots k_{4}}=k_{1}$. Finally, we can perform some explicit numerical checks; figure 1 illustrates that our formula coincides with the result in belzam2009 , in particular, reduces to $p$-deformed volume in the appropriate domain.
@@
  • this is a connected graph with $n$ external “legs”, vertices and the edges connecting some of the vertices. Legs correspond to marked points, vertices — to smooth components of the normalization of degenerate surface and edges to the nodal points where degeneration happens.
  • With each vertex $v$, a number $g_{v}\in\mathbb{Z}_{\geq 0}$ is associated which is a genus of the corresponding smooth component. For each $v$ a stability condition holds: $2g_{v}-2+n_{v}>0$, where $n_{v}$ is the valence of $v$ including edges and legs.
@@
However, the next examples that are presented in the appendix C(g,n)=(0,5)(g,n)=(0,5) and $(g,n)=(1,2)$, although exhibit the expected structure as sum over “stable graphs”, contain additional contributions that do not follow from the diagrammatic technique above.
@@
$\times\left(\prod\limits_{e}\left(\frac{-\partial}{4b^{2}\partial(P_{1,-k_{e}}^{2})}\right)^{n_{e}-1}\prod\limits_{v}\frac{V_{g_{v},n_{v}}^{b}(4\pi bP_{1,-\vec{k}_{v}})}{(2\pi^{2})^{3g_{v}+n_{v}-3}}\right)\Bigg{|}_{P_{e}^{2}=0}$ (57)
The sum over $\Gamma$ includes all stable graphs of genus $g$ with $n$ legs with external markings $k_{1}\dots k_{n}$. For each edge in the graph, a marking $k_{e}$ and a natural number $n_{e}$ are associated and the sum is over all these assignments. It is simpler to write this answer as a sum of tautological intersection numbers on ${\overline{\mathcal{M}}}_{\Gamma}=\prod\limits_{v}{\overline{\mathcal{M}}}_{g_{v},n_{v}}$ — moduli space of stable curves with the degenerations specified by stable graph $\Gamma$, as follows:
@@
$\times\,e^{\sum\limits_{i=1}^{n}(2bP_{1,-k_{i}})^{2}\psi_{i}}\times\sum\limits_{\vec{n_{e}};\,n_{e}=1}^{\infty}\prod\limits_{e}\frac{4^{n_{e}}B_{2n_{e}}(|bP_{1,-k_{e}}|)(-\psi_{e\bullet}-\psi_{e\circ})^{n_{e}-1}}{n_{e}!}$ (58)
In product over $v$ it is understood that each $\kappa$-class is an element in cohomology $H^{\bullet}({\overline{\mathcal{M}}}_{g_{v},n_{v}})$; $\psi_{e\bullet}$ and $\psi_{e\circ}$ are $\psi$-classes, associated to two images of the node point, to which the edge $e$ corresponds, after normalization of the curve. We will not give precise definitions of $\kappa$ and $\psi$-classes and refer to expositions in Zvonkine2012AnIT ; Collier:2023cyw .
@@
It formally reduces to a sum of residues in the right half-plane (at each order in $u$ the series for Hurwitz zeta-function $\zeta(-2n+1,bP)$ is obtained, which reduces to Bernoulli polynomials). This series takes the form
@@
The sine factors appear from Verlinde representation of fusion numbers appearing at the vertices $v_{1},v_{2}$. First, note that this is just a “discrete version” of the “propagator” in CLS Collier:2024matrixmodel , where the integral is replaced by a sum. Second, $P_{1,-n}$ in the total sum go through all the values in Lian-Zuckerman set, enumerating physical operators. The restriction $n\text{ mod }(2p+1)\neq 0$ is in fact not required, since such terms automatically disappear because of sine factors from the fusion constants.
If we substitute the series (59) into (58), we can perform integral over ${\overline{\mathcal{M}}}_{\Gamma}$ explicitly in terms of “quantum volumes”; recall that their definition in our normalization is